# boxedasideAside

### v1

@@ 2.2 The reduced moduli spaces $M_{E,g;n}^{{\rm red}}\subset\overline{M}_{E,g;n}^{{\rm red}}$
Observe that $\pi\circ\varphi$ factors through $[\overline{M}_{E,g}/{\rm Aut}(E)]$, and we already know that the map from $\overline{M}_{E,g}^{red}$ to $[\overline{M}_{E,g}/{\rm Aut}(E)]$ is finite. Therefore, it is enough to show the statement for the map from $[\overline{M}_{E,g}/{\rm Aut}(E)]$ to ¯$\overline{\mathcal{M}}_{g}$. Hence, we want to show that the forgetful map has finite fiber on the smooth locus and generically on each boundary component. Let $g\geq 2$.
  1. 1.
[…]
    Consider a curve $D$ in the image under $\pi\circ\varphi$ of the smooth locus of $\overline{M}^{{\rm red}}_{E,g}$, and let $\eta:D\to E$ be the data of one of the pre-image. We will show that it has only finitely many pre-images in $[M_{E,g}/{\rm Aut}(E)]$. First, note that each double cover $\eta:D\to E$ gives us an element $\sigma\in{\rm Aut}(D)$ which is of order $2$. If we let $\eta^{*}$ and $\sigma^{*}$ to be the induced map on the function field $K(E)$ and $K(D)$ respectively, then we have that $\eta^{*}(K(E))=K(D)^{\sigma^{*}}$. Since $g\geq 2$, we know that ${\rm Aut}(D)$ is a finite group. Therefore, there are finitely many choices for $\sigma$ and hence finitely many choices for $\eta^{*}(K(E))$.
    Now, we show that $\eta^{*}(K(E))$ determines the pre-image. Suppose, we have two pre-images $D\xrightarrow{\eta_{0}}E$ and $D\xrightarrow{\eta_{1}}E$ that satisfy $\eta_{0}^{*}(K(E))=\eta_{1}^{*}(K(E))$. Then there is $\tau\in{\rm Aut}(K(E))$, for which $\eta_{1}^{*}\circ\tau=\eta_{0}^{*}$ (this $\tau$ is obtained from the composition $K(E)\xrightarrow{\eta_{0}^{*}}\eta_{0}^{*}(K(D))\xrightarrow{{\eta_{1}^{*}}^{-1}}K(E)$, which is well defined since $\eta_{0}^{*}(K(E))=\eta_{1}^{*}(K(E))$). Next, there is $f\in{\rm Aut}(E)$ for which $f^{*}=\tau$, so $f\circ\eta_{1}=\eta_{0}$. This shows that $D\xrightarrow{\eta_{0}}E$ and $D\xrightarrow{\eta_{1}}E$ correspond to the same point of $[M_{E,g}/{\rm Aut}(E)]$. We conclude that the forgetful map from $[M_{E,g}/{\rm Aut}(E)]$ to ¯$\overline{\mathcal{M}}_{g}$ is quasi-finite on the smooth locus.
  2. 2.
    Next, consider a boundary component, either $\Xi_{g_{1},g_{2}-1}$ or $\Delta_{g_{1},g_{2}}$. Curves in the generic part of any boundary component are the clutching of two smooth curves, except for one component of $\Xi_{1,g-2}$, when it is the clutching of three smooth curves. Consider a curve $D\in\overline{\mathcal{M}}_{g}$ coming from a generic point of a boundary component. Let $D^{\prime}\xrightarrow{\eta_{0}}C_{0}\xrightarrow{f_{0}}E$ be a preimage of $D$, where $C_{0}$ is the clutching of $E$ with $\mathbb{P}^{1}$ (say $P_{0}$ of $E$ with $Q_{0}$ of $\mathbb{P}^{1}$), and $D$ is the stabilization of $D^{\prime}$ after forgetting the ramification points. Let $D_{a}=\eta_{0}^{-1}(E)$ and $D_{b}=\eta_{0}^{-1}(\mathbb{P}^{1})$, so $g(D_{a})=g_{1}$ and $g(D_{b})=g_{2}$. Then $D^{\prime}$ is the clutching of $D_{a}$ and $D_{b}$, when $\eta_{0}^{-1}(P_{0})$ is identified with $\eta_{0}^{-1}(Q_{0})$. Notice that if $g_{2}\geq 1$, then $D=D^{\prime}$. On the other hand, if $g_{2}=0$, then it must be the case that $(D^{\prime}\to C_{0})\in\Xi_{g-1,0}$. In this case, $D$ is $D_{a}$ with two points in $\eta_{0}^{-1}(P_{0})$ identified.
    1. (a)
      If $g_{1},g_{2}\geq 2$. Then since a curve with genus $\geq 2$ has finite automorphism group, we have only finitely many options for $\sigma_{a}$ and $\sigma_{b}$. We show that $\sigma_{a},\sigma_{b}$ uniquely determine the map $D^{\prime}\to C_{0}\to E$.
      Let $D^{\prime}\xrightarrow{\eta_{1}}C_{1}\xrightarrow{f_{1}}E$ be another preimage of $D$ under $\pi\circ\varphi$. Suppose we have $\eta_{0}^{*}(K(E))=\eta_{1}^{*}(K(E))$ and $\eta_{0}^{*}(K(\mathbb{P}^{1}))=\eta_{1}^{*}(K(\mathbb{P}^{1}))$. Then, there is $h_{a}\in{\rm Aut}(E)$ and $h_{b}\in{\rm Aut}(\mathbb{P}^{1})$ such that $\eta_{1}^{*}\circ h_{a}^{*}=\eta_{0}^{*}$ and $\eta_{1}^{*}\circ h_{b}^{*}=\eta_{0}^{*}$, so $h_{a}\circ\eta_{1}=\eta_{0}\mid_{D_{a}}$ and $h_{b}\circ\eta_{1}=\eta_{0}\mid_{D_{b}}$. Recall that $D^{\prime}$ is the clutching of $D_{a}$ and $D_{b}$, when $\eta_{0}^{-1}(P_{0})\subseteq D_{a}$ is identified with $\eta_{0}^{-1}(Q_{0})\subseteq D_{b}$. Since $D^{\prime}\xrightarrow{\eta_{1}}C_{1}\xrightarrow{f_{1}}E$ is another preimage of $D$, we see that $D$ is also the clutching of $D_{a}$ and $D_{b}$, when $\eta_{1}^{-1}(P_{1})$ is identified with $\eta_{1}^{-1}(Q_{1})$. This means that $\eta_{0}^{-1}(P_{0})=\eta_{1}^{-1}(P_{1})$ and $\eta_{0}^{-1}(Q_{0})=\eta_{1}^{-1}(Q_{1})$. We see that $h_{a}(P_{1})=P_{0}$ and $h_{b}(Q_{1})=Q_{0}$. Thus the map $h=h_{a}\cup h_{b}$ is well defined from $C_{1}$ to $C_{0}$, it is an isomorphism and satisfies $\eta_{0}=h\circ\eta_{1}$.
      Similarly, note that the map $f_{0}:C_{0}\to E$ is identity on $E$ and maps $\mathbb{P}^{1}$ to $P_{0}$, similarly the map $f_{1}:C_{1}\to E$ is identity on $E$ and maps $\mathbb{P}^{1}$ to $P_{1}$. Therefore, we have the following commutative diagram, in which $h$ and $h_{a}$ are isomorphisms:
[…]
      It follows that $D^{\prime}\xrightarrow{\eta_{0}}C_{0}\xrightarrow{f_{0}}E$ and $D^{\prime}\xrightarrow{\eta_{1}}C_{1}\xrightarrow{f_{1}}E$ correspond to the same point of $[\overline{M}_{E,g}/{\rm Aut}(E)]$. Since $\sigma_{a}^{*}$ and $\sigma_{b}^{*}$ determines $\eta_{0}^{*}(K(E))$ and $\eta_{0}^{*}(K(\mathbb{P}^{1}))$, we see that $D$ has only finitely many preimages.
    2. (b)
      If $g_{1}\geq 2$ and $g_{2}=1$. Then $(D^{\prime}\to C_{0})$ must be in $\Delta_{g-1,1}$ or $\Xi_{g-2,1}$. For now we consider both cases simultaneously and we distinguish between the different cases at the end. Note that there are only finitely many options for $\sigma_{a}$. Moreover, $D_{b}$ is an elliptic curve, so its $j$-invariant is the $j$-invariant of the cross-ratio of the four branch points of $\eta_{0}:D_{b}\to\mathbb{P}^{1}$. Thus there are at most six choices for this cross-ratio. We show that $\sigma_{a}$ and the cross ratio uniquely determines $D^{\prime}\to C_{0}\to E$.
      Let $D^{\prime}\xrightarrow{\eta_{1}}C_{1}\xrightarrow{f_{1}}E$ be another preimage of $D$, and suppose that $\eta_{0}^{*}(K(E))=\eta_{1}^{*}(K(E))$ and $\eta_{0}:D_{b}\to\mathbb{P}^{1}$ and $\eta_{1}:D_{b}\to\mathbb{P}^{1}$ have the same cross-ratio of branch points. Then, similar to case (a), there is $h_{a}\in{\rm Aut}(E)$ such that $\eta_{1}^{*}\circ h_{a}^{*}=\eta_{0}^{*}$, so $h_{a}\circ\eta_{1}=\eta_{0}\mid_{D_{a}}$.
      Suppose $\eta_{0}:D_{b}\to\mathbb{P}^{1}$ is ramified at $\{Q_{0},R_{0,1},R_{0,2},R_{0,3}\}$ and $\eta_{1}:D_{b}\to\mathbb{P}^{1}$ is ramified at $\{Q_{1},R_{1,1},R_{1,2},R_{1,3}\}$. Since they have the same cross-ratio, there is a linear fractional transformation $h_{b}:\mathbb{P}^{1}\to\mathbb{P}^{1}$ that sends $\{Q_{1},R_{1,1},R_{1,2},R_{1,3}\}$ to $\{Q_{0},R_{0,1},R_{0,2},R_{0,3}\}$. Now, $\eta_{0}$ and $h_{b}\circ\eta_{1}:D_{b}\to\mathbb{P}^{1}$ have the same ramification points, so $\eta_{0}=h_{b}\circ\eta_{1}$. The rest of the argument is identical to case (a), which shows that $D^{\prime}\xrightarrow{\eta_{0}}C_{0}\xrightarrow{f_{0}}E$ and $D^{\prime}\xrightarrow{\eta_{1}}C_{1}\xrightarrow{f_{1}}E$ correspond to the same point of $[\overline{M}_{E,g}/{\rm Aut}(E)]$.
      Now, when $D$ is in $\Delta_{g-1,1}$, then the point on $D_{b}$ that is clutched with $D_{a}$ is one of the four branch points $\{Q,R_{1},R_{2},R_{3}\}$, so there are finitely many possibilities. On the other hand, when $D$ is in $\Xi_{g-2,1}$, the clutching point of $D_{b}$ is any other point. The position of this point relative to $\{Q,R_{1},R_{2},R_{3}\}=D_{b}[2]$ introduces another dimension to the moduli problem. Two choices of a clutching point give the same point in ¯$\overline{\mathcal{M}}_{g}$ if and only if there is an automorphism of $D_{b}$ that exchanges the two clutching points and preserves $D_{b}[2]$, so again this gives a finite number of preimages in $[\overline{M}_{E,g}/{\rm Aut}(E)]$.
    3. (c)
      If $g_{1}\geq 2$ and $g_{2}=0$. Then, $(D^{\prime}\to C_{0})$ is in $\Xi_{g_{1},0}$, and $D$ is $D_{a}$ with two points in $\eta_{0}^{-1}(P_{0})$ identified with each other. Moreover, $g(D_{b})=0$, so $D_{b}=\mathbb{P}^{1}$. Again $\sigma_{a}$ has only finitely many choices.
      Suppose that $D^{\prime}\xrightarrow{\eta_{1}}C_{1}\xrightarrow{f_{1}}E$ is another preimage of $D$ for which $\eta_{0}^{*}(K(E))=\eta_{1}^{*}(K(E))$. Similar to case (a), there is $h_{a}\in{\rm Aut}(E)$ for which $h_{a}\circ\eta_{1}=\eta_{0}\mid_{D_{a}}$. This forces $h_{a}(P_{1})=P_{0}$. On the other hand, both $\eta_{0}\mid_{D_{b}}$ and $\eta_{1}\mid_{D_{b}}$ are double covers of $\mathbb{P}^{1}$ ramified at two points, so there is $h_{b}\in{\rm Aut}(\mathbb{P}^{1})$ such that $\eta_{0}=h_{b}\circ\eta_{1}$ and $h_{b}(Q_{1})=Q_{0}$. The rest of the proof is the same as in case (a) to show that $D^{\prime}\xrightarrow{\eta_{0}}C_{0}\xrightarrow{f_{0}}E$ and $D^{\prime}\xrightarrow{\eta_{1}}C_{1}\xrightarrow{f_{1}}E$ correspond to the same point of $[\overline{M}_{E,g}/{\rm Aut}(E)]$.
    4. (d)
      If $g_{1}=1$ and $g_{2}\geq 2$. Then $\sigma_{b}$ has only finitely many choices and $\eta_{0}:D_{a}\to E$ has only $4$ choices (it is in $[\overline{M}_{E,1}/{\rm Aut}(E)]$). Suppose that $D^{\prime}\xrightarrow{\eta_{1}}C_{1}\xrightarrow{f_{1}}E$ is another preimage for which $\eta_{0}^{*}(K(\mathbb{P}^{1}))=\eta_{1}^{*}(K(\mathbb{P}^{1}))$ and $\eta_{0},\eta_{1}:D_{a}\to E$ correspond to same point of $[\overline{M}_{E,1}/{\rm Aut}(E)]$. Thus, there is $h_{a}\in{\rm Aut}(E)$ and $h_{b}\in{\rm Aut}(\mathbb{P}^{1})$ such that $\eta_{0}=h_{a}\circ\eta_{1}:D_{a}\to E$ and $\eta_{0}=h_{b}\circ\eta_{1}:D_{b}\to\mathbb{P}^{1}$. The rest of the proof is identical to case (a).
    5. (e)
      If $g_{1}=1$ and $g_{2}=1$, so $D^{\prime}\to E$ is in $\Delta_{1,1}$. This case is identical to case (b), except that $\eta_{0}:D_{a}\to E$ has only $4$ choices as an element of $[\overline{M}_{E,1}/{\rm Aut}(E)]$.
    6. (f)
      If $g_{1}=1$ and $g_{2}=0$, so $D^{\prime}\to E$ is in $\Xi_{1,0}$. This case is identical to case (c), except that $\eta_{0}:D_{a}\to E$ has only $4$ choices as an element of $[\overline{M}_{E,1}/{\rm Aut}(E)]$.
In all cases, we see that $D$ has only finitely many preimages in $[\overline{M}_{E,g}/{\rm Aut}(E)]$. Thus, the map from $[\overline{M}_{E,g}/{\rm Aut}(E)]$ to ¯$\overline{\mathcal{M}}_{g}$ is generically quasi-finite on each boundary divisor. ∎
@@
The preceding proposition is consistent with [SvZ20, Proposition 3.13], which describes the dimension of boundary components of the whole moduli space of Galois $G$-covers.
@@ 3 The $p$-rank stratification of $\mathcal{B}_{E,g}$
Let $E=(E,O_{E})$ be a fixed elliptic curve over an algebraically closed field $k$ of characteristic $p>2$ with $p$-rank $f_{E}$. Consider the locus $\mathcal{B}_{E,g}$ of smooth curves of genus $g\geq 2$ that are double covers of $E$. Our goal is to describe the $p$-rank stratification of $\mathcal{B}_{E,g}$. We achieve this by analyzing its closure in ¯$\overline{\mathcal{M}}_{g}$ using techniques similar to those in [FvdG04], [GP05], and [AP11]. As demonstrated in the proof of Theorem 3.12, the cases when $E$ is ordinary ($f_{E}=1$) and when $E$ is supersingular ($f_{E}=0$) differ significantly, with the latter case requiring particular attention. Consequently, we can describe the $p$-rank stratification of the whole locus $\mathcal{B}_{g}$ of bielliptic curves of genus $g$.
@@
As we have seen in Proposition 2.8, there is a morphism $\overline{M}_{E,g}^{red}\to\overline{\mathcal{M}}_{g}$ that is quasi-finite on the smooth locus and generically quasi-finite on every boundary component. Denote by ¯$\overline{\mathcal{B}}_{E,g}$ the image of this morphism and note that $\mathcal{B}_{E,g}$ coincides with the image of $M_{E,g}^{red}$. Since $\overline{M}_{E,g}^{red}$ is a proper Deligne-Mumford stack by Proposition 2.7 and $M_{E,g}^{red}\subseteq\overline{M}_{E,g}^{red}$ is open and dense, it follows that ¯$\overline{\mathcal{B}}_{E,g}\subseteq\overline{\mathcal{M}}_{g}$ is the closure of $\mathcal{B}_{E,g}\subset\mathcal{M}_{g}$.
Remark 3.1.
@@
Notice that the coarse moduli space of $\mathcal{B}_{E,g}$ is quasi-affine, and consequently, $\mathcal{B}_{E,g}$ cannot contain complete subvarieties. Indeed, let $\widetilde{M}_{E,g}^{\circ}$ denote the open locus of $\widetilde{M}_{E,g}$ where the source curve is smooth. Then $\mathcal{B}_{E,g}$ is the image of $[\widetilde{M}_{E,g}^{\circ}/E]$ under a quasi-finite and proper morphism (hence it induces a finite morphism on the coarse moduli space). The morphism is quasi-finite according to Proposition 2.8, and it is proper because it is the restriction of $[\widetilde{M}_{E,g}/E]\to\overline{\mathcal{B}}_{E,g}$, which are both proper. Hence, it suffices to show that the coarse moduli space of $[\widetilde{M}_{E,g}^{\circ}/E]$ is quasi-affine. If we let $\mathrm{PConf}^{2g-2}(E)$ be the ordered configuration space of $22g-$ 2 points on $E$, then there is a proper étale map $\widetilde{M}^{\circ}_{E,g}\to\mathrm{PConf}^{2g-2}(E)$ that records the branch points of a cover, inducing a finite map on coarse moduli spaces. Now
@@
$[\mathrm{Conf}^{2g-2}(E)/E]\cong[\mathrm{PConf}^{2g-2}(E)/(E\times S_{2g-3})]\cong\mathrm{PConf}^{2g-3}(E-\{O_{E}\})/S_{2g-3}.$
The coarse space of $\mathrm{PConf}^{2g-3}(E-\{O_{E}\})/S_{2g-3}$ is affine because $E-\{O_{E}\}$ is affine and thus the coarse spaces of $[\mathrm{Conf}^{2g-2}(E)/E]$ and $[\widetilde{M}^{\circ}_{E,g}/E]$ are affine.
Remark 3.2.
Also notice that for $g\geq 2$, $\mathcal{B}_{E,g}$ is irreducible. This is because, the map $\overline{M}^{{\rm red}}_{E,g}\to\mathcal{B}_{E,g}$ factors through $[\overline{M}_{E,g}/{\rm Aut}(E)]$ and it was shown in Lemma A.12 that $\overline{M}_{E,g}$ is irreducible.

### v2

@@ 2.2 The reduced moduli spaces $M_{E,g;n}^{{\rm red}}\subset\overline{M}_{E,g;n}^{{\rm red}}$
This can be proved using a careful case-by-case analysis, as is done in the appendix, Proposition A.15. ∎
@@
The preceding proposition is consistent with [SvZ20, Theorem 3.7] and [SvZ20, Proposition 3.13], which describe the dimension of boundary components of the whole locus of Galois $G$-covers in $\overline{\mathcal{M}}_{g}$.
@@ 3 The $p$-rank stratification of $\mathcal{B}_{E,g}$
Let $E=(E,O_{E})$ be a fixed (smooth) elliptic curve over an algebraically closed field $k$ of characteristic $p>2$ with $p$-rank $f_{E}$. Recall that there are two possibilities: $f_{E}=1$ if $E$ is ordinary, and $f_{E}=0$ if $E$ is supersingular. Consider the locus $\mathcal{B}_{E,g}$ of smooth curves of genus $g\geq 2$ that are double covers of $E$. Our goal is to describe the $p$-rank stratification of $\mathcal{B}_{E,g}$. We achieve this by analyzing its closure in $\overline{\mathcal{M}}_{g}$ using techniques similar to those in [FvdG04], [GP05], and [AP11]. As demonstrated in the proof of Theorem 3.12, the cases when $E$ is ordinary and when $E$ is supersingular differ significantly, with the latter case requiring particular attention.
A smooth curve $D$ of genus $g\geq 2$ is called bielliptic if there exists an elliptic curve $E^{\prime}$ and a double cover $D\to E^{\prime}$. We denote by $\mathcal{B}_{g}$ the locus of bielliptic curves in $\mathcal{M}_{g}$. Given an elliptic curve $E^{\prime}$, $\mathcal{B}_{E^{\prime},g}$ can be viewed as a subset of $\mathcal{B}_{g}$. Consequently, we can describe the $p$-rank stratification of the entire locus $\mathcal{B}_{g}$ of bielliptic curves of genus $g$.
@@
As we have seen in Proposition 2.9, there is a morphism $\overline{M}_{E,g}^{{\rm red}}\to\overline{\mathcal{M}}_{g}$ that is quasi-finite on the smooth locus and generically quasi-finite on every boundary component. Denote by $\overline{\mathcal{B}}_{E,g}$ the image of this morphism and note that $\mathcal{B}_{E,g}$ coincides with the image of $M_{E,g}^{{\rm red}}$. Since $\overline{M}_{E,g}^{{\rm red}}$ is a proper Deligne-Mumford stack by Proposition 2.8 and $M_{E,g}^{{\rm red}}\subseteq\overline{M}_{E,g}^{{\rm red}}$ is open and dense, it follows that $\overline{\mathcal{B}}_{E,g}\subseteq\overline{\mathcal{M}}_{g}$ is the closure of $\mathcal{B}_{E,g}\subset\mathcal{M}_{g}$. We summarize some properties of $\mathcal{B}_{E,g}$ and $\overline{\mathcal{B}}_{E,g}$ below.
Lemma 3.1.
For any $g\geq 2$, the moduli space $\overline{\mathcal{B}}_{E,g}$ is irreducible, proper, and has dimension 2g32g-3.
Proof.
Note that $\overline{\mathcal{B}}_{E,g}$ is irreducible for any $g\geq 2$, because the map $\overline{M}^{{\rm red}}_{E,g}\to\overline{\mathcal{B}}_{E,g}$ factors through $\overline{M}_{E,g}/{\rm Aut}(E)]$ and it was shown in Lemma A.13 that $\overline{M}_{E,g}$ is irreducible. The remaining properties follow from Propositions 2.8 and 2.9. ∎
Lemma 3.2.
For any $g\geq 2$, the moduli space $\mathcal{B}_{E,g}$ is irreducible, has dimension 2g32g-3, and does not contain any positive-dimensional complete subvarieties.
Proof.
@@
Lemma 3.1 and Proposition 2.9 imply that $\mathcal{B}_{E,g}$ is irreducible. Moreover, note that the coarse moduli space of $\mathcal{B}_{E,g}$ is quasi-affine, and consequently, $\mathcal{B}_{E,g}$ cannot contain positive-dimensional complete subvarieties. Indeed, let $\widetilde{M}_{E,g}^{\circ}$ denote the open locus of $\widetilde{M}_{E,g}$ where the source curve is smooth. Then $\mathcal{B}_{E,g}$ is the image of $\widetilde{M}_{E,g}^{\circ}/E]$ under a quasi-finite and proper morphism (hence it induces a finite morphism on the coarse moduli space). The morphism is quasi-finite according to Proposition 2.9, and it is proper because it is the restriction of $\widetilde{M}_{E,g}/E]\to\overline{\mathcal{B}}_{E,g}$, which are both proper. Hence, it suffices to show that the coarse moduli space of $\widetilde{M}_{E,g}^{\circ}/E]$ is quasi-affine. If we let $\mathrm{PConf}^{2g-2}(E)$ be the ordered configuration space of 2g22g-2 points on $E$, then there is a proper étale map $\widetilde{M}^{\circ}_{E,g}\to\mathrm{PConf}^{2g-2}(E)$ that records the branch points of a cover, inducing a finite map on coarse moduli spaces. Now
@@
$\mathrm{Conf}^{2g-2}(E)/E]\cong[\mathrm{PConf}^{2g-2}(E)/(E\times S_{2g-3})]\cong\mathrm{PConf}^{2g-3}(E-\{O_{E}\})/S_{2g-3}.$
The coarse space of $\mathrm{PConf}^{2g-3}(E-\{O_{E}\})/S_{2g-3}$ is affine because $E-\{O_{E}\}$ is affine and thus the coarse spaces of $\mathrm{Conf}^{2g-2}(E)/E]$ and $\widetilde{M}^{\circ}_{E,g}/E]$ are affine. ∎